# TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis

### v2

@@ Abstract
Satellite remote sensing from repeated observations and multiple sensors enables a wide range of downstream applications, including climate modeling, carbon accounting, and strategies for conservation and sustainable land use. However, satellite time series are voluminous, often corrupted by sensor noise, clouds, and atmospheric conditions, and unevenly spaced in time, making them challenging to use. We present TESSERA, an open, global, land-oriented remote sensing foundation model that uses self-supervised learning to generate ‘ready-to-use’ embeddings at 10 m scale from pixel-level satellite time series data. TESSERA uses two parallel Transformer-based encoders to combine optical data from ten Sentinel-2 spectral bands at 10-60 m spatial resolution and two Sentinel-1 synthetic aperture radar backscatter coefficients at 10 m resolution to create embeddings that are subsequently fused with a multilayer perceptron to create annual global embedding maps. We compare our work with state-of-the-art task-specific models and other foundation models in five diverse downstream tasks and find that TESSERA closely matches or outperforms these baselines. We believe that TESSERA’s ease of use, openness, computation-, label-, and data-efficiency, and high performance will prove transformative in a wide range of ecological applications.
@@
Monitoring Earth’s dynamic systems through satellite Earth Observation (EO) is critical to addressing global challenges including food security, forest and biodiversity loss, climate change, and disaster mitigation [118, 183, 24]. Petabytes of EO data are available from sensors in various modalities, and declining launch costs promise exponential growth in the future [176]. However, practitioners face a fundamental bottleneck: the scarcity of large, accurately labeled datasets required for supervised training [95]. The direct use of EO data also presents challenges as data quality is affected by sensor noise, cloud cover, atmospheric effects, and nonuniform temporal sampling, preventing full and easy use of this valuable resource [193, 60, 130].
@@
EO data users usually mitigate these issues by annual or seasonal compositing [34], then using supervised machine learning to map from composited images to task-specific classification, such as crop identification [8]. However, compositing removes the critical temporal signal arising from phenology and training a task-specific ML model is computationally onerous. Recent EO approaches tackle the latter problem using foundation models: general-purpose models pre-trained on a broad set of EO data [178]. Their development follows several paths, including supervised pre-training on large labeled datasets for spatio-temporal analysis [21, 20, 160, 18], and various self-supervised learning (SSL) strategies that learn from the inherent structure of the data itself. These SSL techniques include generative methods, such as masked autoencoders [32, 104, 126, 61, 113, 94, 112, 53, 144, 140, 134, 67, 149], diffusion-based models [192, 80, 115], and contrastive methods that learn efficient representations by comparing data samples [14, 102, 7, 99, 81, 23, 59, 70, 169]. Subsequently, these general purpose models are adapted and fine-tuned for specific remote sensing tasks such as high-resolution (semantic) segmentation [161, 177, 184, 28, 191].
@@
Although existing SSL approaches create rich multidimensional ‘embeddings’ from unlabeled data through auxiliary tasks that reveal latent patterns [15, 97], they still do not address the loss of the temporal signal during compositing. In recent work, Lisaius et al [95] have demonstrated that the Barlow Twins loss [187], grounded in redundancy reduction principles [17], can be adapted to preserve the spectral-temporal signal in latent embeddings when processing dense EO time series. Building on this insight, we introduce TESSERA, a pixel-oriented foundation model for EO time series that creates 128-dimensional latent embeddings at 10m resolution and global scale to enable state-of-the-art performance across a diverse array of complex tasks. Our experiments demonstrate that these TESSERA-derived embeddings outperform baseline methods in tasks as diverse as estimating canopy height in rainforests in Borneo, detecting crop types in Austria, and identifying fire scars in California.
@@
Importantly, our outputs are ‘ready to use’: Users need neither to fine-tune the model, nor download and process voluminous satellite data, nor perform inferencing other than what is needed to perform a downstream task. This makes TESSERA frugal in terms of both data handline and computation. They can also keep their data private instead of having to upload it to a centralized repository. This removes a crucial barrier to use. We demonstrate that TESSERA is label-efficient: Only a relatively small number of labels for task-specific training are needed to achieve state-of-the-art performance. Moreover, our model is open source, enabling extensibility and scientific reproducibility.
@@
We have made a gap-free, global, annual, 10-meter resolution embedding map for 2024 openly available through the Geotessera Python library [152] and aim to eventually provide global maps for the entire period 2017-2024. These information-rich precomputed maps greatly simplify a wide range of downstream EO applications. We have also released interactive and batch tools for habitat mapping using TESSERA embeddings (see [1] for details).
@@ 1 Overview of the TESSERA System
TESSERA uses self-supervised learning to create per-pixel embeddings from the unlabeled time-series of observations from Sentinel-1 (SAR) and Sentinel-2 Multispectral Instruments (MSI), learning complementary information from the optical and microwave data streams. It comprises two main components: a dual-branch encoder and a projector network (see Fig. 1 for details).
@@
TESSERA models its input as a set of ‘d-pixels.’ Each d-pixel represents the annual time series of Sentinel-1 and Sentinel-2 observations at a specific 10 m pixel on Earth as a two-dimensional array ordered in time (that is, channels $\times$ timesteps). This format preserves the temporal evolution of spectral/backscatter signatures inherent in the satellite data. Our design accommodates common observational gaps and irregularities, such as those due to cloud cover or orbit overlap, by masking invalid rows, which are handled during subsequent processing stages. Model training and inferencing is done by sparse temporal sampling followed by flattening of these d-pixels (see Appendix A.4 and Fig. 1 for details). For each 10 m pixel, the fused TESSERA encoders generate a compact (8 bit), 128-dimensional embedding that encapsulates that pixel’s annual temporal and spectral characteristics. Under typical mid-latitude observation conditions, this compresses input data by a factor of $\approx$ 24, with higher compression ratios at higher latitudes.
@@
Figure 1: Architecture and data processing pipeline of the TESSERA foundation model. a, Overview of the end-to-end workflow. Spatially aligned time-series data from Sentinel-2 MSI and Sentinel-1 SAR are first extracted for each pixel location, forming modality-specific ‘d-pixels’. For self-supervised training, two distinct augmentations are created from each d-pixel–each augmentation is a random cloud-free sample from an annual time series of Sentinel-2 observations along with the corresponding Sentinel-1 data–processed through a dual-branch architecture with weights shared between the branches. The resulting 128-dimensional embeddings from both modalities are fused by a multilayer perceptron (MLP) to create the output embeddings of the system. The fused features are then expanded to 16,384 dimensions by a large projector network. A modified Barlow Twins loss ($\mathcal{L}_{BT}+\mathcal{L}_{MIX}$) is calculated on the cross-correlation matrix of these projected features, optimizing the model to learn augmentation-invariant and decorrelated embeddings that are internally decorrelated, and fairly invariant to obscured days in the Sentinel-2 time series. b, Detailed structure of a single encoder branch. A d-pixel, represented as a matrix of temporal steps versus spectral bands (40 time-steps $\times$ 2 bands for Sentinel-1), is first embedded. A Day-of-Year-based temporal encoding is added to these embeddings. The resulting sequence is then processed by an 4-block Transformer Encoder. Finally, a Gated Recurrent Unit (GRU) pooling layer processes the sequence of temporal features; its final hidden state, which summarizes the entire time series, serves as the single 128-dimensional embedding for that modality, which is then passed to the fusion MLP shown in panel a.
@@
TESSERA uses two separate, parallel Transformer-based encoders: one for Sentinel-1 SAR backscatter data in the C-band (VV and VH polarizations) and another for Sentinel-2 MSI data (10 spectral bands covering the visible, near, and shortwave infrared). Each encoder processes the temporal sequence of a d-pixel, utilizing multi-head self-attention mechanisms to capture complex temporal dependencies and patterns specific to each data modality. Positional encodings derived from the Day-of-Year (DoY) are incorporated to provide temporal context. An attention pooling layer within each encoder aggregates the temporal features into a fixed-size vector that represents the annual signature for that modality [1]. See Appendix A.7 and Figure 1 for architectural details.
@@
Modality-specific embeddings derived from dual encoders are first fused using a multilayer perceptron (MLP) which is trained alongside the encoders. This fused embedding, which serves as the 128-dimensional pixel embedding for downstream tasks, is then quantized before being fed into a high-dimensional projector network. This projector, a significantly larger MLP, expands the dimensionality of the fused representation to facilitate effective redundancy reduction during the self-supervised learning phase. A detailed breakdown of the model’s architectural parameters is provided in Appendix A.6 (Table 2).
@@ 1.1 Model Training
TESSERA learns the mapping of inputs to embeddings through a self-supervised pretraining process on a large-scale, unlabeled global dataset. It is trained by optimizing a modified Barlow Twins loss function [187, 95]. The core principle is to learn embeddings that are invariant to differing views of the same pixel’s optical-band data (e.g., different subsets of cloud-free observation dates) while minimizing redundancy between the learned feature dimensions. This is achieved by generating two distinct augmented views for each input d-pixel, through independent sparse temporal samples. This avoids creating augmentations by applying artificial (e.g. hand-crafted) distortions. The model then learns to reconcile these partial views. The loss function includes the standard Barlow Twins terms for invariance and redundancy reduction, complemented with an additional mix-up regularization term to enhance robustness and intersample interaction during training (see Appendix A.7.2 for the full loss function, Equation 11). In practice, we randomly sub-sample each d-pixel three times, to create a total of three pairs of augmentations.
The pretraining phase utilizes a vast global dataset comprising approximately 800 million d-pixels, down-sampled from Sentinel-1 and Sentinel-2 imagery (we use a downsampling ratio of 400:1). To ensure successful generalization, a multilevel data sampling, shuffling, and batching strategy was developed. This ensured that each training mini-batch contained a highly diverse set of d-pixels from various geographic locations and acquisition conditions, preventing model overfitting and promoting the learning of universally applicable features. The model was trained for one epoch on this dataset using distributed training techniques (see Appendix A.7.3 for further details on data preparation, the shuffling strategy and optimization). The key hyperparameters used during pretraining are summarized in Appendix A.7.3 (Table 3).
@@ 1.2 Model Inference and Global Embeddings Maps
Following pretraining, the TESSERA dual encoder (with frozen weights and excluding the projector) is used to generate an annual 128-dimensional embedding for every 10-meter pixel globally. This involves making inferences from a sample of length 40 from the annual Sentinel-1 and Sentinel-2 time series for each pixel through the trained encoders. To ensure that our maps are gap-free, random sampling with replacement is used to generate augmentations from d-pixels that have fewer than 40 non-cloudy observations. The primary output of this inference stage is a set of annual, global, 10-meter resolution TESSERA embedding maps. These maps are designed as readily usable multichannel geospatial data layers. This “Embeddings-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, precomputed features can be directly ingested by downstream models without the need for raw satellite data processing or running the TESSERA model itself (see Appendix A.8). We are generating global embeddings for 2024 (approx. 250 TB) and will progressively extend coverage backward year by year until 2017.
@@ 2 Downstream Tasks
The effectiveness and generalizability of the TESSERA embeddings are evaluated on a diverse range of downstream remote sensing tasks. These include, but are not limited to, pixel-wise classification (e.g., crop type mapping), pixel-wise regression (e.g., canopy height and AGB estimation) and patch-based dense predictions (for example, semantic segmentation). In these evaluations, the pre-trained TESSERA encoder acts as a fixed feature extractor; given adequate computing resources, fine-tuning of the model is also possible. Lightweight, task-specific model heads (e.g., shallow MLPs or UNet-style decoders) are then trained using these fixed embeddings. This approach typically requires substantially less labelled data and computational resources compared to training deep models from scratch, yet achieves or surpasses the state of the art. See Appendix D and Supplementary Fig. 10 for the general downstream application workflow.
@@ 2.1 Crop Type Classification
To evaluate performance, we used the INVEKOS Austrian crop dataset [3] from the 2021–2022 growing season. This is an extensive data set that covers 1850 km2 east of Vienna with 154 different crop types, which we consolidated into 17 classes based on crop similarity (e.g., merging feed beet with sugar beets) [3]. We compared the performance of a lightweight classification head trained on TESSERA embeddings against three key baselines: (1) an RF classifier trained on engineered temporal features from raw satellite data, (2) the PRESTO foundation model [149], and (3) Google Satellite Embedding (GSE) [56]. For the foundation models, we applied the same lightweight head to their respective embeddings for a direct comparison. We evaluated performance using multiple metrics, including F1 scores, under varying data availability scenarios, from 30% of labelled data to one-shot learning. We also evaluated performance on a patch-based semantic segmentation task against several other foundation models. Details on data processing, model architectures, and training protocols are provided in Appendix E.
TESSERA consistently surpasses both baselines in all training regimes, especially in settings where labeled data are scarce (Fig. 2). In pixel-wise classification tasks with training data splits from 1% to 30%, a simple MLP trained on TESSERA embeddings achieves higher weighted and macro F1 scores than an identical MLP trained on PRESTO or GSE embeddings, and substantially outperforms the RF baseline (Fig. 2a). TESSERA also maintains statistically significant advantages in one-shot and few-shot scenarios (Fig. 2b). Furthermore, TESSERA’s utility extends to patch-based semantic segmentation, where it consistently achieves higher mIoU and macro F1 scores compared to other leading foundation models, including PRESTO and GSE, across various patch sizes (Fig. 2c).
These performance gains are due to the high quality of the learned embeddings. A UMAP analysis [103] of the embedding space reveals that TESSERA embeddings form tighter, more semantically coherent clusters by crop class compared to those from both PRESTO and GSE (Fig. 2d). This improved separability highlights the strength of TESSERA as a foundational tool for agricultural applications.

### v3

@@ Abstract
Satellite remote sensing from repeated observations and multiple sensors enables a wide range of downstream applications, including habitat mapping, carbon accounting, and strategies for conservation and sustainable land use. However, satellite time series are voluminous, often corrupted by sensor noise, clouds, and atmospheric conditions, and unevenly spaced in time, making them challenging to use. We present TESSERA, an open, global, land-oriented remote sensing foundation model that uses self-supervised learning to generate ‘ready-to-use’ embeddings at 10 m scale from pixel-level satellite time-series data. TESSERA uses two parallel Transformer-based encoders to combine optical data from ten Sentinel-2 spectral bands at 10-20 m spatial resolution and two Sentinel-1 synthetic aperture radar backscatter coefficients at 10 m resolution to create embeddings that are subsequently fused with a multilayer perceptron to create annual global embedding maps. We compare our work with state-of-the-art task-specific models and other foundation models in five diverse downstream tasks and find that TESSERA closely matches or outperforms these baselines. We believe that TESSERA’s ease of use, state-of-the-art performance, openness, and computation- and labelled data-efficiency will prove transformative in a wide range of ecological applications.
@@
Monitoring Earth’s dynamic systems through satellite Earth Observation (EO) is critical to addressing global challenges including food security, forest and biodiversity loss, climate change, and disaster mitigation [117, 185, 25]. Petabytes of EO data are available from sensors in various modalities, and declining launch costs promise continued growth in the future [178]. However, practitioners face a fundamental bottleneck: the scarcity of large, accurately labelled datasets required for supervised training [95]. The direct use of EO data also presents challenges as data quality is affected by sensor noise, cloud cover, atmospheric effects, and non-uniform temporal sampling, preventing full and easy use of this valuable resource [196, 61, 131].
@@
EO data users usually mitigate these issues by annual or seasonal compositing [35], then using supervised machine learning to map from composited images to task-specific classification, such as crop identification [9]. However, compositing removes the critical temporal signal arising from phenology (i.e. timing of biological events such as flowering) and training a task-specific ML model is computationally onerous. Recent EO approaches tackle the latter problem using foundation models: general-purpose models pre-trained on a broad set of EO data [180]. These include both supervised pre-training on large labelled datasets for spatio-temporal analysis [22, 21, 162, 19], and self-supervised learning (SSL) strategies that learn from the inherent structure of the data itself. Common SSL techniques include generative methods, such as masked autoencoders [33, 136, 68, 151], diffusion-based models  [195, 81, 114], and contrastive methods [15, 60, 71, 171]. General purpose foundation models must be adapted (or ‘fine-tuned’) for specific downstream tasks such as habitat mapping and semantic segmentation [163, 179, 186, 29, 193].
@@
Although existing SSL approaches create rich multidimensional ‘embeddings’ from unlabelled data through auxiliary tasks that reveal latent patterns [16, 97], they still do not address the loss of the temporal signal during compositing. In recent work, Lisaius et al. [95] have demonstrated that the Barlow Twins loss [189], grounded in redundancy reduction principles [18], can be adapted to preserve the spectral-temporal signal in latent embeddings when processing dense EO time series. Building on this insight, we introduce TESSERA, a pixel-oriented foundation model for EO time series that creates 128-dimensional latent embeddings at 10 m resolution and global scale to enable state-of-the-art performance across a diverse array of complex tasks. Our experiments demonstrate that these TESSERA-derived embeddings outperform state-of-the-art methods in tasks as diverse as estimating canopy height in rainforests in Borneo, detecting crop types in Austria, mapping above-ground biomass in Finland, and identifying fire scars in California.
@@
Importantly, our outputs are ‘ready to use’: users need neither to fine-tune the model, nor download and process voluminous satellite data, nor perform inferencing other than what is needed to perform a downstream task. This makes TESSERA frugal in terms of both data handling and computation. They can also keep their data private instead of having to upload it to a centralized repository. This removes a crucial barrier to use. We demonstrate that TESSERA is label-efficient: Only a relatively small number of labels for task-specific training is needed to achieve state-of-the-art performance. Moreover, our model is open source, enabling extensibility and scientific reproducibility.
@@
We have made a gap-free, global, annual, 10 m resolution embedding map for 2024 openly available through the Geotessera Python library [154] and aim to eventually provide global maps for the entire period 2017-2024. These information-rich precomputed maps greatly simplify a wide range of downstream EO applications. We have also released interactive and batch tools for habitat mapping using TESSERA embeddings (see [1] for details).
@@ 1 Overview of the TESSERA System
TESSERA uses self-supervised learning to create per-pixel embeddings from a global dataset of unlabelled time-series of observations from Sentinel-1 Synthetic Aperture Radar (SAR) and Sentinel-2 Multispectral Instruments (MSI), learning complementary information from the optical and microwave data streams. It is trained by optimizing a modified Barlow Twins loss function [189]. The original Barlow Twins approach seeks to create embeddings that are a) invariant to augmentations of its input, where these augmentations are typically rotations, reflections, and cropping of an image and b) whose dimensions are as decorrelated as possible. Inspired by the work by Lisaius et al. [95], we use the Barlow Twins approach to learn embeddings that are invariant to the choice of random subsets of cloud-free observations in the Sentinel-2 optical time series, fused with the corresponding Sentinel-1 data, while minimizing redundancy between the learned feature dimensions. This is achieved by randomly sampling non-cloudy days from the Sentinel-2 annual observation time series so that model then learns to reconcile these partial views, interpolating missing data.
The loss function includes the standard Barlow Twins terms for invariance and redundancy reduction, complemented with an additional mix-up regularization term to enhance robustness and inter-sample interaction during training (see Appendix A.7.2 for details, Equation 11). We now discuss the system architecture.
@@
TESSERA comprises two main components: a dual-branch encoder and a projector network (see Fig. 1 for details). TESSERA models its input as a set of ‘d-pixels.’ Each d-pixel represents the annual time series of Sentinel-1 and Sentinel-2 observations at a specific 10 m pixel on Earth as a two-dimensional array ordered in time (that is, channels $\times$ timesteps). This format preserves the temporal evolution of the spectral/backscatter signatures inherent in the satellite data, while allowing observational gaps and irregularities, such as those due to cloud cover or orbit overlap. Rows that are cloud-corrupted or otherwise invalid are indicated by a binary mask and ignored when creating augmentations. Model training and inferencing is done by random temporal sampling followed by flattening of these d-pixels (see Appendix A.4 and Fig. 1 for details). In practice, we randomly sub-sample each d-pixel three times, to create a total of three pairs of augmentations. For each 10 m pixel, the fused TESSERA encoders generate a compact (8 bit), 128-dimensional embedding that encapsulates that pixel’s annual temporal and spectral characteristics. Under typical mid-latitude observation conditions, this compresses the input data by a factor of $\approx$ 24, with higher compression ratios at higher latitudes.
@@
Figure 1: Architecture and data processing pipeline of the TESSERA foundation model. a, Overview of the end-to-end workflow. Spatially aligned time-series data from Sentinel-2 MSI and Sentinel-1 SAR observations are first extracted for each pixel location, to form mode-specific ‘d-pixels’ (shown as vertical rectangle, with cloud-corrupted rows indicated in gray). To allow self-supervised training, two augmentations are created from each d-pixel. Each augmentation is a cloud-free random sample from the corresponding d-pixel. Augmentations are processed through a dual-branch architecture with weights shared between the branches, which is the key feature of Barlow Twin SSL. The two resulting 128-dimensional embeddings are fused by a multilayer perceptron (MLP) to create the final output embeddings. During training (but not inferencing) the fused features are expanded to 16,384 dimensions by a large projector network. A modified Barlow Twins loss (a linear combination of $\mathcal{L}_{BT}$ and $\mathcal{L}_{MIX}$) is calculated from the cross-correlation matrix of these projected features, forcing the model to learn dimensionally decorrelated embeddings that are invariant to masked (i.e, cloud-corrupted) days in the Sentinel-2 d-pixel. k b, Detailed structure of a single encoder branch. A d-pixel, represented as a matrix of temporal steps versus spectral bands (40 time-steps $\times$ 2 bands for Sentinel-1), is first embedded using a single-layer MLP (not shown). A Day-of-Year-based temporal encoding is added to this initial embedding. The resulting sequence is processed by an 4-block Transformer Encoder. Finally, a Gated Recurrent Unit (GRU) pooling layer processes the sequence of temporal features; its final hidden state, which summarizes the entire time series, serves as the single 128-dimensional embedding for that modality, which is then passed to the fusion MLP shown in panel a.
@@
TESSERA uses two separate, parallel Transformer-based encoders: one for Sentinel-1 SAR backscatter data in the C-band (VV and VH polarizations) and another for Sentinel-2 MSI data (10 spectral bands covering the visible, near, and shortwave infrared). Each encoder processes a d-pixel, utilizing multi-head self-attention mechanisms to capture the temporal dependencies and patterns specific to each data modality. Positional encodings derived from the Day-of-Year (DoY) are incorporated to provide temporal context. An attention pooling layer within each encoder aggregates the temporal features into a fixed-size vector that represents the annual signature for that modality [1]. See Appendix A.7 and Figure 1 for architectural details.
@@
Modality-specific embeddings derived from dual encoders are fused using a multilayer perceptron (MLP) which is trained alongside the encoders. This fused embedding, which serves as the 128-dimensional pixel embedding for downstream tasks, is then quantized before being fed into a high-dimensional projector network. This projector, a significantly larger MLP, expands the dimensionality of the fused representation to facilitate redundancy reduction during the self-supervised learning phase. A detailed breakdown of the model’s architectural parameters is provided in Appendix A.6 (Table 2).
@@ 1.1 Model Training
The pretraining phase utilizes a large global dataset of approximately 800 million d-pixels, down-sampled 400:1 from Sentinel-1 and Sentinel-2 time series. To ensure successful generalization, a multilevel data sampling, shuffling, and batching strategy was developed. This ensured that each training mini-batch contained a highly diverse set of d-pixels from various geographic locations and acquisition conditions, preventing model overfitting and promoting the learning of universally applicable features. The model was trained for one epoch in this dataset using distributed training techniques (see Appendix A.7.3 for further details on data preparation, the shuffling strategy, and optimization). The key hyperparameters used during pretraining are summarized in Appendix A.7.3 (Table 3).
@@ 1.2 Model Inference and Global Embeddings Maps
Following pretraining, the TESSERA dual encoder (with frozen weights and excluding the projector) is used to generate an annual 128-dimensional embedding for every 10 m pixel globally. This involves making inferences from a sample of length 40 from the annual Sentinel-1 and Sentinel-2 time series for each pixel through the trained encoders. To ensure that our maps are gap-free, random sampling with replacement is used to generate augmentations from d-pixels that have fewer than 40 non-cloudy observations. The primary output of this inference stage is a set of annual, global, 10 m resolution TESSERA embedding maps. These maps are designed as readily usable multichannel geospatial data layers. This “Embeddings-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, precomputed features can be directly ingested by downstream models without the need for raw satellite data processing or running the TESSERA model itself (see Appendix A.8). We are generating global embeddings for 2024 (approx. 250 TB) and will progressively extend coverage backward year by year until 2017. Visualizing the first three dimensions of the embeddings as R,G, and B channels reveals rich, detailed textures corresponding to diverse land cover and land use patterns. These visualizations highlight the model’s ability to capture information across multiple scales, from individual agricultural fields to entire continents (Fig. 2).
N
Figure 2: TESSERA embeddings capture fine-grained geospatial patterns across multiple scales. The first three dimensions of the 128-dimensional embeddings are mapped to RGB channels for visualization. a, Examples from diverse global locations, showing a variety of landforms and vegetation patterns. b, A multi-scale view of a complex agricultural landscape. b.1 shows the overall region, with subsequent panels (b.2, b.3) zooming into the white-boxed areas, revealing progressively finer details down to individual crop fields. c, Continental-scale visualizations demonstrating the richness and consistency of the embeddings over vast areas, including (c.1) the Continental United States, (c.2) Europe, and (c.3) the Amazon basin.
@@ 2 Downstream Tasks
The effectiveness and generalizability of the TESSERA embeddings are evaluated on a diverse range of downstream remote sensing tasks. These include pixel-wise classification, such as for crop type mapping, pixel-wise regression, such as for estimating canopy height and carbon density of forests, and patch-based dense predictions, such as for semantic segmentation. (See Appendix D and Supplementary Fig. 12 for the general downstream application workflow.) In these evaluations, the pre-trained TESSERA encoder, with frozen weights, acts as a feature extractor (although, given adequate computing resources, fine-tuning the model is also possible), converting d-pixels from a region of interest to embeddings. Lightweight, task-specific model heads (e.g., shallow MLPs or UNet-style decoders) are then trained using these embeddings, acting as explanatory variables in regression modelling. This approach is computation- and labelled-data efficient: it typically requires substantially less labelled data and computational resources compared to training deep models from scratch, yet achieves or surpasses the state-of-the-art.
@@ 2.1 Crop Type Classification
To evaluate performance, we used the INVEKOS Austrian crop dataset [3] from the 2021–2022 growing season. This is an extensive data set that covers 1850 km2 east of Vienna with 154 different crop types, which we consolidated into 17 classes based on crop similarity (e.g., merging feed beet with sugar beets) [3]. We compared the performance of a small classifier neural network trained using TESSERA embeddings with three key state-of-the-art alternatives: (1) an RF classifier trained on engineered temporal features from raw satellite data, (2) the PRESTO foundation model [151], and (3) Google Satellite Embedding (GSE) [57]. For the foundation models, we used the same small neural network to classify their respective embeddings to allow a direct comparison. We evaluated performance using multiple metrics, including F1 scores, under varying data availability scenarios, from 30% of labelled data to one-shot learning. We also evaluated performance on a patch-based semantic segmentation task against several other foundation models. Details on data processing, model architectures, and training protocols are provided in Appendix E.
TESSERA consistently surpasses both baselines in all training regimes, especially in settings where labelled data are scarce (Fig. 3). In pixel-wise classification tasks with training data splits from 1% to 30%, a simple MLP trained on TESSERA embeddings achieves higher weighted and macro F1 scores than an identical MLP trained on PRESTO or GSE embeddings, and substantially outperforms the RF baseline (Fig. 3a). TESSERA also maintains statistically significant advantages in one-shot and few-shot scenarios (Fig. 3b). Furthermore, TESSERA’s utility extends to patch-based semantic segmentation, where it consistently achieves higher mIoU and macro F1 scores compared to other leading foundation models, including PRESTO and GSE, across various patch sizes (Fig. 3c).
These performance gains are due to the high quality of the learned embeddings. A UMAP analysis [102] of the embedding space reveals that TESSERA embeddings form tighter, more semantically coherent clusters by crop class compared to those from PRESTO and GSE (Fig. 3d). This improved separability highlights the strength of TESSERA as a foundational tool for agricultural applications.